OpenStudy (anonymous):

lim(2e^x sinx) As x is approaching pi/2?

OpenStudy (anonymous):

Is this the product of 2e^x sin(x)?

OpenStudy (anonymous):

@trogdortheburninator I'm not sure what the product of means.. It says to solve by substitution, so I would assume it's (2e^pi/2)*sin(pi/2), I'm just not sure as to what the e means in the equation.

OpenStudy (anonymous):

use wolframalpha

OpenStudy (anonymous):

then you should be set

OpenStudy (anonymous):

\[\huge \lim_{x \rightarrow \frac{\pi}{2}}(2e^x\sin x)\] \[\huge =\lim_{x \rightarrow \frac{\pi}{2}} Use substitution for e^x.

OpenStudy (anonymous):

e is a number like pi which is used for integration and also for applying this number for other applications of calculus.

OpenStudy (anonymous):

Sorry, different question!

OpenStudy (anonymous):

By looking at it, I can see that sin(x) should go to 1, and e^x should go to whatever e^2pi is